OpenStudy (anonymous):

Simplify square root of 81x^13 (assume that all variales are non-negative real numbers

OpenStudy (campbell_st):

\[\sqrt{81x^{13}} = \sqrt{9^2\times(x^6)^2\times x} = \sqrt{9^2}\times \sqrt{(x^6)^2}\times \sqrt{x}\] gives \[9x^6\sqrt{x}\]